# Wiedemann-Franz law violation domain for graphene and nonrelativistic systems

### v2

@@ Abstract
Systematic and comparative research on Lorenz ratios for graphene and nonrelativistic systems has been studied to identify their Wiedemann-Franz law violation domain. Fermi energy and temperature are the main governing parameters for deciding the values of the Lorenz ratio, which is basically thermal conductivity divided by electrical conductivity times temperature times Lorenz number. Metals as three-dimensional nonrelativistic electron gas locate at higher Fermi energy by temperature domain, where Lorenz ratio remains one. Hence, they obey the Wiedemann-Franz law. By creating higher doping in a two-dimensional graphene system, one can again reach a higher Fermi energy by temperature domain and get a constant Lorenz ratio. For both graphene and nonrelativistic systems, the Lorenz ratio goes below one if we go lower Fermi energy by temperature domain, which is possible for the graphene system by decreasing the doping concentration. Experimentally observed greater than one Lorenz ratio in this lower Fermi energy by temperature domain or Dirac fluid domain indicates that non-fluid expressions of Lorenz ratio should be replaced by fluid-type expressions. We have noticed a divergent trend of Lorenz ratio in the Dirac fluid domain using its fluid-type expression, and it matches with the trend of experimental data.
@@ I Introduction
where $L_{0}$ is known as the Lorenz number. The fact is very well known from our text book Refs. [1, 2, 3, 4, 5]. In natural unit, $\hbar=c=k_{B}=1$ and $e^{2}=\frac{4\pi}{137}$, we can express temperature $T$ in eV, $\sigma$ in eV and $\kappa$ in eV2. So Lorenz number will be a dimensionless ratio $L_{0}=\frac{137\pi}{12}\approx 35.84$. In the present paper, we will mostly use the natural unit methodology for convenience, but we may go to the actual unit in some cases (whenever necessary).
@@
The Wiedemann-Franz (WF) law has proven remarkably robust for many metallic systems, where electrons are the main electric and thermal charge transportation carriers. Due to the similar mechanism of transportation in free electron theory [1, 2, 3, 4, 5], the dimensionless (in natural unit) ratio of two transport coefficients - thermal and electrical conductivity becomes constant. However, deviations have been observed in many systems like $Li_{0.9}Mo_{6}O_{17}$ [6], $CeCoIn_{5}$ [7], $ZrZn_{2}$ [8], $YbRh_{2}Si_{2}$ [9], $(Pr,Ce)_{2}CuO_{4}$ [10] and $VO_{2}$ [11]. For our convenience, if we define the ratio of thermal and electrical conductivity as $L=\frac{\kappa}{\sigma T}$, and if we call $L/L_{0}$ as Lorenz ratio, then the validity of the WF law will be established from the relations $L=L_{0}$ or $\frac{L}{L_{0}}=1$. On the other hand, relations $\frac{L}{L_{0}}>1$ or $\frac{L}{L_{0}}<1$ reflect violation of the WF law. A large diverging outcomes $\frac{L}{L_{0}}>10^{4}$ is expected in one-dimensional exotic Tomonaga-Luttinger liquids [6]. On the other hand, a strong downward violation of the WF Law $L<L_{0}$ was observed in $WP_{2}$ semimetal [12, 13] and MoP binary compound [14]. This downward violation is also observed in heavy fermion metals [7], marginal ferromagnetic metal [8], anti-ferromagnetic metal [9] and copper-oxide superconducting material [10] at low-temperature regions as well as in metallic vanadium dioxide [11] at high-temperature range ($240$-$340^{\circ}$). Understanding the WF law violation mechanism of that wide range of systems is a non-trivial task and challenge for the theoretical community to explain.
@@
Recently, Crossno et al. [15] have found a similar kind of the WF law violation in graphene systems by tuning the doping concentration and temperature. Unlike standard metal, the Fermi level of graphene can be moved upwards or downwards from the relative Dirac point depending upon the n- or p-type doping of graphene. Fermi energy $(\epsilon_{F})$ or a chemical potential $(\mu)$ at Dirac point is considered zero for undoped graphene and, via n- or p-type doping, will shift it towards the positive or negative direction. The net charge carrier concentration will be tuned from zero to non-zero when one experimentally goes from undoped to doped graphene cases, and theoretically, one goes from $\epsilon_{F}=0$ to $\epsilon_{F}\neq 0$ cases. This kind of tuning possibility of $\epsilon_{F}$ can not be expected for metal systems as it almost remains constant. Their typical values remain within the range $\epsilon_{F}=2-10$ eV, for which one can expect the limit $\epsilon_{F}/T>>1$ at room temperature $T\approx 0.025$ eV. In this limit, one can consider electrons in metal as Fermi gas and use Sommerfeld expansion, which provides a linear T dependent on electron-specific heat. Fermi gas is completely based on the crude non-interacting assumption, but there is a theory of interacting Fermions system, which is popularly known as the Fermi Liquid (FL) Theory. Originally, Landau [16, 17, 18] proposed this phenomenological theory for studying 3He. In FL prescription, interaction is taken care of via the effective mass of electrons by assuming a quasi-particle picture. Hence, some mass-independent quantities like $L$ remain the same as $L_{0}$ for both FL and Fermi gas prescriptions. So, if we define $\epsilon_{F}/T>1$ as the FL domain and $\epsilon_{F}/T>>1$ as the Fermi gas domain, then electron-doped graphene systems almost follow WF law. When we go towards undoped or clean graphene systems with $\epsilon_{F}/T<1$, Fermi liquid theory becomes invalid (as it is commonly accepted). It is concluded from the experimental observation of WF law violation in this $\epsilon_{F}/T<1$ or $\epsilon_{F}/T<<1$ domains, popularly called Dirac fluid (DF) or Dirac Liquid (DL) domain. Theoretical works [19, 20, 21, 22, 23] are attempted to explain this WF law violation in the DF domain. In this regards, electron hydrodynamics (eHD) or fluid dynamics in graphene system, recently observed by Refs [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], is thought to be linked with this WF law violation. In this context, the present work is planned to explore a transition from non-fluid to fluid type expressions of Lorenz ratio for a 2-dimensional (2D) graphene (G) system with linear energy-momentum relation ($E\propto p$) as well as a 3-dimensional (3D) metal system with nonrelativistic (NR) energy-momentum relation ($E\propto p^{2}$). For mathematical completeness, we have demonstrated all possible cases like 2D-NR, 3D-NR, 2D-G, and 3D-G but our final destination is kept to show how fluid expressions of 2D-G case can be associated with WF law violation and how non-fluid expressions of 3D-NR case, followed by metal systems, obey WF law. In the direction of eHD, our earlier comparative research on ideal [38] and viscous dissipating [39] parts of energy-momentum tensor calculation clearly demonstrate that the graphene hydrodynamics is neither relativistic hydrodynamics nor nonrelativistic hydrodynamics. The present work has gone through similar comparative research on electrical conductivity, thermal conductivity, and Lorenz ratio for a systematic search of the WF law violation domain.
@@
The article is organized as follows, in Sec. (II), the formalism part of WF law calculations for different cases. The results and discussions are in Sec. (III) and summarised by Sec.( IV).
@@ II Formalism
Graphene is a 2-dimensional single atomic layer of carbon atoms tightly bound into a honeycomb lattice [40]. Near the Dirac point, electrons in graphene follow the linear dispersion relation
@@
where $v_{F}$ is the Fermi velocity of electrons in graphene, whose values remain within the range $0.003c$-$0.01c$ [41]. In one direction, electrons in graphene do not follow a quadratic relation between energy and momentum ($\epsilon=p^{2}/(2m)$) like in the nonrelativistic case. On the other hand, its velocity ($v_{F}$) is not close to the speed of light ($c$), so no relativistic corrections are expected. In that sense, electrons in graphene follow neither relativistic nor nonrelativistic. So, its different properties may not be the same as expected for traditional nonrelativistic or relativistic matters as shown in Refs. [38, 39], for thermodynamics [38] and dissipative viscous [39] properties. In this regard, thermal and electrical conductivity for the graphene (G) case may also be interesting to compare with nonrelativistic (NR) and ultrarelativistic (UR) cases, which are studied in the present work. In sub-sec. (II.1), we have gone through the first non-fluid type framework description, which is traditionally adopted in standard books of solid state physics [1, 2, 3, 4, 5]. Next, in sub-sec. (II.2), we have extended our framework towards fluid type description.
@@ II.1 Non-fluid description
For systematic study and understanding, we will calculate the Lorenz ratios for the 3D system and then the 2D system. The NR, UR, and G case expressions of the Lorenz ratio will be addressed step by step for each system. Then, we will also examine the limiting conditions using Sommerfeld expansion.
3D-NR: Let us first discuss the 3D-NR case. The well-known Drude’s formula for the electrical conductivity of nonrelativistic electrons in 3D solids (metals) is
$\sigma_{NR}^{3D}=\frac{ne^{2}\tau_{c}}{m}=\frac{ne^{2}\lambda}{mv_{F}},$ (3)
and the thermal conductivity is
$\kappa_{NR}^{3D}=\frac{1}{3}nv_{F}\lambda\,\prescript{3D}{NR}{[C_{V}]_{e}},$ (4)
where, $n$ is the number (electron) density, $\tau_{c}$ is the relaxation time, $m$ is electron mass, $\lambda=v_{F}\tau_{c}$ is the mean free path, and $[C_{V}]_{e}$ is the electronic specific heat per particle. After taking the ratio of thermal and electrical conductivity, we get
where $\epsilon_{F}=\mu=\frac{1}{2}mv_{F}^{2}$ is the Fermi energy.
Here, in Eq. (5), $[C_{V}]_{e}$ is the main important thing to calculate the conductivity ratio. We will use the two definitions of $[C_{V}]_{e}$:
  1. 1.
    $[C_{V}]_{e1}=\frac{\partial}{\partial T}\Big{(}\frac{U}{N}\Big{)}\Bigg{|}_{V,\ \epsilon_{F}}$,
  2. 2.
    $[C_{V}]_{e2}=\frac{1}{N}\frac{\partial U}{\partial T}\Bigg{|}_{V,\ \epsilon_{F}}$.
Here, for subsequent simplicity, we will use the notations like NR1 and G1 concerning $[C_{V}]_{e1}$ and NR2 and G2 dealing with $[C_{V}]_{e2}$. The former definition prescribes taking $T$ derivative of internal energy per particle $u=U/N$, while the latter definition says to take $T$ derivative of internal energy $U$ and then normalized by $N$. The detailed calculation can be seen in Appendix C. Let us first put the former specific heat,
[$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]},$ (6)
in Eq. (5), the Lorenz ratio for 3D-NR will be
Here,
[…]
is the standard Fermi integral function (see details in Appendix B) with $A=e^{\epsilon_{F}/k_{B}T}$. Using Sommerfeld lemma, in the limit of $\epsilon_{F}/k_{B}T=lnA>>1$, the electronic specific heat becomes
[$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{\pi^{2}}{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (9)
and then Eq. (7) becomes
$L_{NR1}^{3D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (10)
which is the so-called Wiedemann-Franz law for metals.
Next, using another definition of $[C_{V}]_{e2}$ for the 3D-NR case, whose general expression will be (see Appendix C)
[$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}~{},$ (11)
we will get the Lorenz ratio as
[$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3\pi^{2}}{4}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (13)
and then the Eq. (12) becomes
$L_{NR2}^{3D}=\frac{\pi^{2}}{2}\left(\frac{k_{B}}{e}\right)^{2}=1.5L_{0}.$ (14)
Reader can see that Eq. (10) and (14), both are independent of $T$ , $\epsilon_{F}$ and are made by universal constants like $k_{B}$ and $e$. So, the domain of $T-\epsilon_{F}$ plane, where $\frac{L}{L_{0}}$ or $L$ becomes independent of $T$ , $\epsilon_{F}$ , can be considered as Wiedemann-Franz law obeying domain and remaining will be Wiedemann-Franz law violating domain. Although, in terms of $L_{0}$ , $L$ is $1.5$ factor times larger for Eq. (14).
and the thermal conductivity,
$\kappa_{G}^{3D}=\frac{1}{3}nv_{F}\lambda\prescript{3D}{G}{[C_{V}]_{e}},$ (16)
will form the ratio
[$\prescript{3D}{G1}{[C_{V}]_{e1}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]}$ (18)
@@
[$\prescript{3D}{G2}{[C_{V}]_{e2}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (19)
@@
and
[$\prescript{3D}{G1}{[C_{V}]_{e1}}=\pi^{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (22)
and
[$\prescript{3D}{G2}{[C_{V}]_{e2}}=3\pi^{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (23)
and so, Eqs. (20) and (21) become
$L_{G1}^{3D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (24)
and
$L_{G2}^{3D}=\pi^{2}\left(\frac{k_{B}}{e}\right)^{2}=3L_{0},$ (25)
respectively.
[$\prescript{2D}{NR1}{[C_{V}]_{e1}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]},$ (27)
and
[$\prescript{2D}{NR2}{[C_{V}]_{e2}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (28)
in Eq. (26), we get the Lorenz ratios
and
$L_{NR1}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (31)
$L_{NR2}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0}.$ (32)
2D-G: Now, we are taking an actual 2D graphene case. The ratio of $\kappa_{G}^{2D}$ and $\sigma_{G}^{2D}$ is given by
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]},$ (34)
and
[$\prescript{2D}{G2}{[C_{V}]_{e2}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (35)
in Eq. (33), we get the Lorenz ratios
$\frac{L_{G1}^{2D}}{L_{0}}=\frac{\kappa_{G1}^{2D}}{\sigma_{G1}^{2D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}$ (36)

### v3

@@ Abstract
A systematic non-fluid to fluid transition framework and comparative research on Lorenz ratios for graphene and nonrelativistic systems have been studied to identify their Wiedemann-Franz law violation domain. Here, Lorenz ratio is defined as thermal conductivity divided by electrical conductivity times temperature times Lorenz number. In non-fluid framework, Lorenz ratio become exactly one, which means that the Wiedemann-Franz is obeyed within a Fermi Liquid domain. When one enters from Fermi Liquid to Dirac Fluid domain, Lorenz ratio becomes less than one in non-fluid framework but in fluid framework, it always remain greater than one for both domain. By compiling our outcomes and connecting with experimental data, a non-fluid to fluid transition framework is expected during the transition from Fermi Liquid to Dirac Fluid domain.
@@ I Introduction
where $L_{0}$ is known as the Lorenz number. The fact is very well known from our text book Refs. [ashcroft2022solid, devanathan2021wiedemann, Pathria:1996hda, pillai2006solid, puri1997solid]. In natural unit, $\hbar=c=k_{B}=1$ and $e^{2}=\frac{4\pi}{137}$, we can express temperature $T$ in eV, $\sigma$ in eV and $\kappa$ in eV2. So Lorenz number will be a dimensionless ratio $L_{0}=\frac{137\pi}{12}\approx 35.84$. In the present paper, we will mostly use the natural unit methodology for convenience, but we may go to the actual unit in some cases (whenever necessary).
@@
The Wiedemann-Franz (WF) law has proven remarkably robust for many metallic systems, where electrons are the main electric and thermal charge transportation carriers. Due to the similar mechanism of transportation in free electron theory [ashcroft2022solid, devanathan2021wiedemann, Pathria:1996hda, pillai2006solid, puri1997solid], the dimensionless (in natural unit) ratio of two transport coefficients - thermal and electrical conductivity becomes constant. However, deviations have been observed in many systems like $Li_{0.9}Mo_{6}O_{17}$ [wakeham2011gross], $CeCoIn_{5}$ [tanatar2007anisotropic], $ZrZn_{2}$ [smith2008marginal], $YbRh_{2}Si_{2}$ [pfau2012thermal], $(Pr,Ce)_{2}CuO_{4}$ [hill2001breakdown] and $VO_{2}$ [lee2017anomalously]. For our convenience, if we define the ratio of thermal and electrical conductivity as $L=\frac{\kappa}{\sigma T}$, and if we call $L/L_{0}$ as Lorenz ratio, then the validity of the WF law will be established from the relations $L=L_{0}$ or $\frac{L}{L_{0}}=1$. On the other hand, relations $\frac{L}{L_{0}}>1$ or $\frac{L}{L_{0}}<1$ reflect violation of the WF law. A large diverging outcome $\frac{L}{L_{0}}>10^{4}$ is expected in one-dimensional exotic Tomonaga-Luttinger liquids [wakeham2011gross]. On the other hand, a strong downward violation of the WF Law $L<L_{0}$ was observed in $WP_{2}$ semimetal [WP2_1_departure, WP2_2_departure] and MoP binary compound [MoP_departure]. This downward violation is also observed in heavy fermion metals [tanatar2007anisotropic], marginal ferromagnetic metal [smith2008marginal], anti-ferromagnetic metal [pfau2012thermal] and copper-oxide superconducting material [hill2001breakdown] at low-temperature regions as well as in metallic vanadium dioxide [lee2017anomalously] at high-temperature range ($240$-$340^{\circ}$K). Similar violation is also observed for the nuclear matter, produced in heavy ion collision experiments [VDWHRG_2023, CSPM_2019, HRG1_2019, HRG2_2023]. Understanding the WF law violation mechanism of that wide range of systems is a non-trivial task and challenge for the theoretical community to explain.
@@
Recently, Crossno et al. [crossno2016observation] have found a similar kind of the WF law violation in graphene systems by tuning the doping concentration and temperature. Unlike standard metals, the Fermi level of graphene can be moved upwards or downwards from the relative Dirac point depending upon the n- or p-type doping of graphene. Fermi energy $(\epsilon_{F})$ or a chemical potential $(\mu)$ at Dirac point is considered zero for undoped graphene and, via n- or p-type doping, will shift it towards the positive or negative direction. The net charge carrier concentration will be tuned from zero to non-zero when one experimentally goes from undoped to doped graphene cases, and theoretically, one goes from $\epsilon_{F}=0$ to $\epsilon_{F}\neq 0$ cases. This kind of tuning possibility of $\epsilon_{F}$ can not be expected for metal systems as it almost remains constant. Their typical values remain within the range $\epsilon_{F}=2-10$ eV, for which one can expect the limit $\epsilon_{F}/T>>1$ at room temperature $T\approx 0.025$ eV. In this limit, one can consider electrons in metal as Fermi gas and use Sommerfeld expansion, which provides a linear T dependent on electron-specific heat. Fermi gas is completely based on the crude non-interacting assumption, but there is a theory of interacting Fermions system, which is popularly known as the Fermi Liquid (FL) Theory[gochan2020fermi]. Originally, Landau [Landau1, Landau2, Landau3] proposed this phenomenological theory for studying 3He. In FL prescription, interaction is taken care of via the effective mass of electrons by assuming a quasi-particle picture. So, if we define $\epsilon_{F}/T>1$ as the FL domain and $\epsilon_{F}/T>>1$ as the Fermi gas (FG) domain, then electron-doped graphene systems almost follow WF law. When we go towards undoped or clean graphene systems with $\epsilon_{F}/T<1$, Fermi liquid theory becomes invalid (as it is commonly accepted). It is concluded from the experimental observation of WF law violation in this $\epsilon_{F}/T<1$ or $\epsilon_{F}/T<<1$ domains, popularly called Dirac fluid (DF) or Dirac Liquid (DL) domain. Theoretical works [rycerz2021wiedemann, tu_Yin_23_the, Theory_G_WF, Mahajan:2013cja, Lucas:2018kwo, zarenia2019breakdown, PhysRevLett.115.056603, PhysRevB.99.085104] are attempted to explain this WF law violation in the DF domain. In this regards, electron hydrodynamics (eHD) or fluid dynamics in graphene system, recently observed by Refs [exp1, exp2, PRB_21, PRB_2_21, PRB_3_21, PRR, Nat_Com_21, PRB_coll, Nat_19, Sci_19, Hall_Sci_19, Nat_Tec_19, Nat_Com_18, Sci_16_Res], is thought to be linked with this WF law violation. In the direction of material perspective [principi2015violation], one can link eHD and WF law violation as follows. The electron-electron scattering length becomes larger than the electron-impurity or electron-phonon scatterings in DF domain but reverse is occured in FL or Fermi gas domain. We know that the relaxation time is inverse to scattering strength. So, the e-e scattering dominating domain i.e. DF domain shows eHD behaviour, where electron relaxation time becomes much much smaller than the system size. In this context, the present work is planned to explore a transition from non-fluid to fluid type expressions of Lorenz ratio for a 2-dimensional (2D) graphene (G) system with linear energy-momentum relation ($E\propto p$) as well as a 3-dimensional (3D) metal system with nonrelativistic (NR) energy-momentum relation ($E\propto p^{2}$). For mathematical completeness, we have demonstrated all possible cases like 2D-NR, 3D-NR, 2D-G, and 3D-G but our final destination is kept to show how fluid expressions of 2D-G case can be associated with WF law violation and how non-fluid expressions of 3D-NR case, followed by metal systems, obey WF law. In the direction of eHD, our earlier comparative research on ideal [TZW] and viscous dissipating [CWA] parts of energy-momentum tensor calculation clearly demonstrate that the graphene hydrodynamics is neither relativistic hydrodynamics nor nonrelativistic hydrodynamics. The present work has gone through similar comparative research on electrical conductivity, thermal conductivity, and Lorenz ratio for a systematic search of the WF law violation domain.
@@
The article is organized as follows, in Sec. II, the formalism part of WF law calculations for different cases. The results and brief interpretation are in Sec. III and Sec. IV respectively. At the end, our study is concluded in Sec. V.
@@ II Formalism
Graphene is a 2-dimensional single atomic layer of carbon atoms tightly bound into a honeycomb lattice [geim2007rise]. Near the Dirac point, electrons in graphene follow the linear dispersion relation
@@
where $v_{F}$ is the Fermi velocity of electrons in graphene, whose values remain within the range 0.003c0.003c-$0.01c$ [vg_nature]. In one direction, electrons in graphene do not follow a quadratic relation between energy and momentum ($\epsilon=p^{2}/(2m)$) like in the nonrelativistic case. On the other hand, its velocity ($v_{F}$) is not close to the speed of light ($c$), so no relativistic corrections are expected. In that sense, electrons in graphene follow neither relativistic nor nonrelativistic. So, its different properties may not be the same as expected for traditional nonrelativistic or relativistic matters as shown in Refs. [TZW, CWA], for thermodynamics [TZW] and dissipative viscous [CWA] properties. In this regard, thermal and electrical conductivity for the graphene (G) case may also be interesting to compare with nonrelativistic (NR) and ultrarelativistic (UR) cases.
@@ II.1 Non-fluid description
For systematic study and understanding, we will calculate the Lorenz ratios for the 3D system and the 2D system. The expressions for 2D graphene are derived and analyzed in this section, highlighting its distinct transport characteristics. Meanwhile, the corresponding calculations for other systems, including 3D non-relativistic (NR), 3D graphene (G), and ultra-relativistic (UR) cases in both 3D and 2D, are provided in Appendix D for completeness. Then, we will also examine the limiting conditions using Sommerfeld expansion.
2D-G: For actual 2D graphene case, the electrical and thermal transport properties are:
$\sigma_{G}^{2D}=\frac{ne^{2}\tau_{c}v_{g}^{2}}{\mu}~,$ (3)
$\sigma_{Q}^{3D}=\frac{ne^{2}\tau_{c}c^{2}}{\mu_{q}}~,$ (4)
$\kappa_{G}^{2D}=\frac{1}{2}nv_{F}^{2}\tau_{c}\prescript{2D}{G}{[C_{V}]_{e}}~,$ (5)
the ratio of $\kappa_{G}^{2D}$ and $\sigma_{G}^{2D}$ is given by
$\frac{\kappa_{G}^{2D}}{\sigma_{G}^{2D}}=\frac{1}{2}\frac{\epsilon_{F}}{e^{2}}\prescript{2D}{G}{[C_{V}]}.$ (7)
the ratio of $\kappa_{Q}^{3D}$ and $\sigma_{Q}^{3D}$ is given by
$\frac{\kappa_{Q}^{3D}}{\sigma_{Q}^{3D}}=\frac{1}{3}\frac{\mu_{q}}{e^{2}}\prescript{3D}{Q}{[C_{V}]_{e}}.$ (8)
$\frac{\kappa_{Q}^{3D}}{\sigma_{Q}^{3D}}=\frac{1}{3}\frac{\mu_{q}}{e^{2}}\prescript{3D}{Q}{[C_{V}]}.$ (9)
$\prescript{2D}{G1}{[C_{V}]_{e}}=2k_{B}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg],$ (10)
@@
$\prescript{2D}{G2}{[C_{V}]_{e}}=2k_{B}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg],$ (11)
@@
$\frac{L_{G1}^{2D}}{L_{0}}=\frac{\kappa_{G}^{2D}}{\sigma_{G}^{2D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg]$ (12)